At a luncheon for doctors, all but 40 were neurosurgeons, all but 50 were pediatricians, and all but 60 were cardiologists. How many of each group attended the luncheon?
Answer by TimothyLamb(4379)

---
N = total number of doctors:
x = N - 40
y = N - 50
z = N - 60
---
y + z = 40
x + z = 50
x + y = 60
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x= neurosurgeons = 35
y= pediatricians = 25
z= cardiologists = 15
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php